OpenStudy (anonymous):

estimate using compatible number 1.29+4.69+3.45+1.59

OpenStudy (anonymous):

12

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

1.3 + 4.7 + 3.5 + 1.6 = 11.1

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

Rounded to the nearest 1/10th The actual sum is shown below: 1.29 + 4.69 + 3.45 + 1.59 = 11.02

OpenStudy (anonymous):

Compatible Numbers? Never heard of such an estimation procedure until this problem. 6+5 = 11 ?

OpenStudy (anonymous):

its so hard to explain

OpenStudy (anonymous):

I read the your yahoo site. My view is that to line up or manipulate the numbers so that they add up to factors of ten requires more mental energy than to just do the calculation bye hand. Do what you have to do to pass the class. In the real world you would use a hand held calculator.

OpenStudy (anonymous):

im sure it is not to hard

OpenStudy (anonymous):

its only grade 8 math

OpenStudy (anonymous):

the trouble is that i dont really understand the technique

OpenStudy (anonymous):

it would be 11

OpenStudy (anonymous):

if you're supposed to get a whole number.